# Quantum Kernel Methods under Scrutiny: A Benchmarking Study

### v2

@@ Abstract
Since the entry of kernel theory in the field of quantum machine learning, quantum kernel methods (QKMs) have gained increasing attention with regard to both probing promising applications and delivering intriguing research insights. Two common approaches for computing the underlying Gram matrix have emerged: fidelity quantum kernels (FQKs) and projected quantum kernels (PQKs). Benchmarking these methods is crucial to gain robust insights and to understand their practical utility. In this work, we present a comprehensive large-scale study examining QKMs based on FQKs and PQKs across a manifold of design choices. Our investigation encompasses both classification and regression tasks for five dataset families and $64$ datasets, systematically comparing the use of FQKs and PQKs quantum support vector machines and kernel ridge regression. This resulted in over $20,000$ models that were trained and optimized using a state-of-the-art hyperparameter search to ensure robust and comprehensive insights. We delve into the importance of hyperparameters on model performance scores and support our findings through rigorous correlation analyses. In this, we also closely inspect two data encoding strategies. Moreover, we provide an in-depth analysis addressing the design freedom of PQKs and explore the underlying principles responsible for learning. Our goal is not to identify the best-performing model for a specific task but to uncover the mechanisms that lead to effective QKMs and reveal universal patterns.
I Introduction
@@
Within the rapidly evolving field of quantum machine learning (QML) [1, 2], quantum kernel methods (QKMs) [3, 4, 5, 6, 7, 8, 9] have emerged as a particularly interesting and promising branch of research. For example, Liu et al. [10] have proven a rigorous quantum advantage for a classification task engineered from the discrete logarithm problem, although it requires a fault-tolerant quantum computer due to its reliance on a Shor-type data encoding. Beyond that, Refs. [3, 11] formally reveal that supervised QML models are kernel methods. In particular this implies that QKMs can be embedded into the rich mathematical framework of conventional kernel theory [12]. Among other appealing properties, this guarantees that when training a kernel-based model one finds the optimal solution due to the convexity of the training landscape.
@@
The key idea of kernel methods are kernel functions that implicitly map input data into a higher-dimensional space where the learning problem becomes linear. Quantum kernel methods leverage the principles of quantum mechanics to perform these mappings into the exponentially large Hilbert space of quantum states. This is done using unitary operators, with input data encoded into rotation angles. In practice, this is realized by parameterized quantum circuits (PQCs) [3, 13, 14, 8], which in the context of quantum kernels are referred to as data encoding circuits. Two common approaches to evaluate the corresponding quantum kernel functions have become established: fidelity quantum kernels (FQKs) [3, 4, 15, 7] and projected quantum kernels (PQKs) [16, 9, 17]. A unified framework for generalized trace-induced quantum kernels that includes FQKs and linear PQKs has been proposed recently in Ref. [9].
@@
Wide-ranging application-oriented research in diverse domains emerged from the promising potentials of QKMs. This ranges from financial classification tasks [18] and quantum support vector machines (QSVM) for modeling transition probabilities in health and disability insurance [19] over a quantum kernel classifier for real high-dimensional data taken from the domain of cosmology [5] to QKMs for solving differential equations [20]. Apart from that, several works aims to understand the impact of certain properties of the underlying data encoding circuits on the model performance. In this regard, Hubregtsen et al. [21] performed a numerical study on the relation between classification accuracy and the PQC’s expressibility and entangling capability. Similar works [6, 22] explore the effect of entanglement in data encoding circuits on the performance of different classification tasks using FQKs. The importance of data preprocessing in FQKs is demonstrated in Ref. [23] based on experiments with real-world classification datasets. Moreover, there are recent efforts to automize the encoding circuit architecture search, cf., e.g., Ref. [24] and references therein.
@@
Despite the variety of previous work, a comprehensive understanding of the wealth of design choices and their underlying mechanisms for various datasets is still incomplete. Our work attempts to make a contribution to closing this gap further by systematically analyzing the diversity of design criteria on a large scale and systematically examining both FQK and PQK approaches. We address this through state-of-the-art hyperparameter search and correlation analysis. A schematic view on the scope of this work is shown in Fig. 1.
@@
We conduct an extensive benchmarking study for classification and regression tasks using various QKMs such as quantum support vector classification (QSVC), quantum kernel ridge regression (QKRR), and quantum support vector regression (QSVR). Additionally, we generate a broad and general database by exploring nine popular data encoding circuits from the literature, offering a diverse analysis on the impact of the choice of encoding circuit. Beyond that, we propose two different encoding strategies to introduce feature redundancies and systematically compare them. Furthermore, the present work delves deep into the different design aspects of PQKs by examining the effects of measurement operators and outer kernels on model performance. We additionally address the question of which components of the PQK definition are responsible for learning, i.e., the projected quantum circuits versus the outer kernel. Finally, we numerically support all findings by rigorous correlation analysis. In total, this study resulted in over $20,000$ quantum kernel models that were trained to provide an extensive database and to ensure robust and comprehensive insights. To facilitate this, we developed a software tool. Its basic functionality is indicated in Fig. 1.
@@
Previous studies have focused mostly on classification tasks. For example, Bowles et al. [34] conducted a large-scale evaluation of popular QML models, including Quantum Neural Networks (QNNs), FQKs and PQKs across various datasets. While QKMs where part of the study, the influence of the different design choices was not a particular focus. Their work was preceded by first attempts to systematically explore certain aspects of quantum model design, cf., e.g., Refs [35, 36]. Egginger et al. [37] have conducted a hyperparameter study for QKMs which extends the findings of Ref. [29] for PQKs. However, unlike this work, they exclusively focus on a single feature map (Hamiltonian evolution). In contrast to these previous findings, the present study offers an in-depth analysis of QKMs that significantly broadens the scope of investigation in multiple aspects and aims to answer aspects that have been previously untouched. This includes addressing regression problems, encoding mechanisms and a detailed analysis of PQKs. By extending previous findings we aim to contribute to a holistic understanding of QKMs.
@@
This work is organized as follows. Section II introduces the theoretical basics of QKMs. Hereafter, we give detailed insights about the study design, including an overview on the multitude of different quantum kernel models considered in this work, an introduction of the datasets as well as implementation details for the corresponding hyperparameter optimization pipeline and the final experimental setup. In Sec. IV we first thoroughly investigate model performances and the influence of hyperparameters for classification and regression tasks across all datasets, encoding circuits and quantum kernel models and support corresponding findings with correlation analyses. Secondly, we carefully examine different data encoding strategies and give an in-depth analysis on different design options within the PQK approach, i.e., choice of outer kernel function and measurement operator. We discuss universal findings and patterns across all experiments and comment on the necessity of entanglement in data encoding circuits in Sec. V.
II Theoretical Background
@@
One of the most interesting aspects of QKMs is that they can be formally embedded into the rich and powerful mathematical framework of classical kernel theory [3, 11]. The key idea behind the conventional kernelized approach to (supervised) machine learning is to find and analyze patterns by transforming the respective learning problem from the original input data domain $\mathcal{X}$ to a higher-dimensional (potentially infinite-dimensional) feature space $\mathcal{F}$, where the learning tasks attains a trivial form, i.e., becomes a linear model. This mapping is accomplished by a feature map $\phi:\mathcal{X}\to\mathcal{F};\mathbf{x}\mapsto\phi(\mathbf{x})$. Kernels, are real- or complex-valued symmetric and positive semi-definite functions of two input data points, $k:\mathcal{X}\times\mathcal{X}\to\mathbb{K}$, where $\mathbb{K}$ can be $\mathbb{R}$ or $\mathbb{C}$, respectively. In this regard, another central concept is that of reproducing kernel Hilbert space (RKHS) $\mathcal{H}_{k}$, which is an alternative feature space of a kernel; more precisely a feature space of functions constructed from the kernel. The RKHS uniquely determines the kernel and vice versa [38] and in addition, for every kernel there exists at least one feature map such that [11]
@@
$h(\mathbf{x})$ $=\sum_{i=1}^{N}c_{i}k(\mathbf{x},\mathbf{x}_{i})\,.$ (2)
Note that the number of terms in the sum is independent of the dimension of $\mathcal{H}_{k}$. The determination of the coefficients $\{c_{i}\}$ is a convex optimization problem [40]. This result is central to kernel methods used for supervised learning problems, e.g., SVMs [12], Gaussian Processes [41] or KRR [42]. All these algorithms are based on the kernel Gram matrix $G_{ij}=[k(\mathbf{x}_{i},\mathbf{x}_{j})]$.
In QML we process input data by encoding (embedding) them into quantum states of the form
|$\ket{\psi_{\bm{\theta}}(\mathbf{x})}$ $=U(\mathbf{x},\bm{\theta})\ket{0}^{\otimes n}\,,$ (3)
where $\bm{\theta}$ are variationally trainable parameters to refine the embedding into the quantum Hilbert space $\mathcal{H}^{Q}$. In practice, the unitary encoding operator $U(\mathbf{x},\bm{\theta})$ can be implemented by a data encoding quantum circuit which manipulates an initial $n$-qubit quantum state |$\ket{0}^{\otimes n}$. This reveals the striking similarity to kernel methods: Both utilize mathematical frameworks that map information into high-dimensional spaces for processing. In particular, as shown in Refs. [3, 11], the central concept of QKMs is that they can be formulated as a classical kernel method (e.g. SVM or KRR) whose kernel is computed using a quantum computer. Since quantum computations inherently feature the quantum mechanical principles of superposition and entanglement, the resulting quantum kernels hold the prospect of designing machine learning models that are able to learn complex problems that are out of reach for conventional machine learning methods [10].
@@
Figure 2: Schematic illustration of the basic working principle of QKMs and its two most common approaches to compute respective quantum kernel Gram matrices. Data points are mapped from the input space $\mathcal{X}$ to the quantum Hilbert space $\mathcal{H}^{Q}$ by encoding them into quantum states |$\ket{\psi(\mathbf{x},\bm{\theta})}$. Access to $\mathcal{H}^{Q}$ is provided by measurements, which can be expressed by inner products of quantum states in full analogy to classical kernel theory. Left: By using the native geometry of the quantum state space, i.e., the Hilbert-Schmidt inner product and leveraging this fidelity-type metric to define quantum kernels leads to FQKs, cf. Eq. (4). Right: Instead of directly processing quantum states within the quantum Hilbert space it has been shown that it can be beneficial to first project them to an approximate classical representation using, e.g., reduced physical observables. This concept gives rise to the family of PQKs. One of the simplest forms of defining PQKs is given in Eq. (8) and corresponds to measuring $k$-particle reduced density matrices and process the result with a classical kernel function $\kappa$. In both cases (FQK and PQK) the resulting kernel Gram matrices are subsequently passed to a classical kernel algorithm.
To formalize this, we define the density matrix $\rho_{\bm{\theta}}(\mathbf{x})=\outerproduct{\psi(\mathbf{x},\bm{\theta})}{\psi(\mathbf{x},\bm{\theta})}$ | as the corresponding data-encoding feature map [11, 4]. With this, we can leverages the native geometry of the quantum state space and naturally define a quantum kernel using the Hilbert-Schmidt inner product, i.e.,
$k^{\text{FQK}}_{\bm{\theta}}(\mathbf{x},\mathbf{x}^{\prime})$ $=\mathrm{tr}\left[\rho_{\bm{\theta}}(\mathbf{x})\rho_{\bm{\theta}}(\mathbf{x}^{\prime})\right]\,.$ (4)
@@
$k^{\text{FQK}}_{\bm{\theta}}(\mathbf{x},\mathbf{x}^{\prime})$ $=\left|\Braket{\psi_{\bm{\theta}}(\mathbf{x})}{\psi_{\bm{\theta}}(\mathbf{x}^{\prime})}\right|^{2}\,,$ (5)
which, since it represents a fidelity-type metric, is referred to as the fidelity quantum kernel (FQK) [16]. Computing the full Gram matrix according to Eq. (5), in practice requires to execute two circuits and to sample the expectation value of the projector $P_{0}=\ket{0}^{\otimes n}\bra{0}^{\otimes n}$, of a $n$-qubit system [8].
@@
$\rho_{\bm{\theta}}^{K}(\mathbf{x})$ $=\mathrm{tr}_{j\notin K}\left[\rho_{\bm{\theta}}(\mathbf{x})\right]\,,$ (6)
@@
$f_{k}(\mathbf{x},\bm{\theta})$ $=\mathrm{tr}(\rho_{\bm{\theta}}^{K}(\mathbf{x})O)\,,$
$=\mathrm{tr}(\rho_{\bm{\theta}}(\mathbf{x})O^{k})\,,$ (7)
where $O^{k}$ represents a $k$-local measurement operator acting on $k\leq n$ qubits, e.g. $P^{\otimes k}\otimes\mathds{1}^{\otimes(n-k)}$, with $P$ a Pauli operator. As such, PQKs can be generally defined as
$k^{\text{PQK}}_{\bm{\theta}}(\mathbf{x},\mathbf{x}^{\prime})$ $=\kappa\left(f_{k}[\mathbf{x},\bm{\theta}],f_{k}[\mathbf{x}^{\prime},\bm{\theta}]\right)\,.$ (8)
@@
$k^{\text{PQK}}_{\bm{\theta}}(\mathbf{x},\mathbf{x}^{\prime})$ $=\exp(-\gamma\sum_{k,P}\left[\mathrm{tr}\{P\rho_{\bm{\theta}}^{k}(\mathbf{x}\bm{)}\}-\mathrm{tr}\{P\rho_{\bm{\theta}}^{k}(\mathbf{x}^{\prime})\}\right]^{2})$
$=\exp\left[-\gamma F_{\bm{\theta}}(\mathbf{x},\mathbf{x}^{\prime})\right]\,,$ (9)
where $\gamma\in\mathbb{R_{+}}$ is a hyperparameter and $\rho_{\bm{\theta}}^{k}(\mathbf{x})$ is the 1-RDM, i.e., the partial trace of the quantum state $\rho_{\bm{\theta}}(\mathbf{x})$ over all qubits except for the $k$-th qubit; cf. Eq. (6). Here, we introduced the latter notation for later use in Sec. IV.5. Note that unless otherwise stated, we are referring to the form as given in Eq. (9) when considering PQKs.
III Study Design
In this study, we systematically investigate the interplay of various hyperparameters in QKMs with the aim to identify patterns and mechanisms that enhance model performance (see Fig. 1). The following section details the methodological aspects of our study design, explaining the models and datasets we consider as well as providing insight into the implementation to realize this comprehensive analysis.
III.1 Models

### v3

@@ Abstract
Since the entry of kernel theory in the field of quantum machine learning, quantum kernel methods (QKMs) have gained increasing attention with regard to both probing promising applications and delivering intriguing research insights. Benchmarking these methods is crucial to gain robust insights and to understand their practical utility. In this work, we present a comprehensive large-scale study examining QKMs based on fidelity quantum kernels (FQKs) and projected quantum kernels (PQKs) across a manifold of design choices. Our investigation encompasses both classification and regression tasks for five dataset families and 64 datasets, systematically comparing the use of FQKs and PQKs quantum support vector machines and kernel ridge regression. This resulted in over 20,000 models that were trained and optimized using a state-of-the-art hyperparameter search to ensure robust and comprehensive insights. We delve into the importance of hyperparameters on model performance scores and support our findings through rigorous correlation analyses. Additionally, we provide an in-depth analysis addressing the design freedom of PQKs and explore the underlying principles responsible for learning. Our goal is not to identify the best-performing model for a specific task but to uncover the mechanisms that lead to effective QKMs and reveal universal patterns.
keywords:
quantum computing, quantum machine learning, quantum kernel methods, fidelity quantum kernels, projected quantum kernels, benchmarking, hyperparameter optimization
1 Introduction
@@
Within the rapidly evolving field of quantum machine learning (QML) [6, 12], quantum kernel methods (QKMs) [50, 21, 42, 56, 24, 27, 18] have emerged as a particularly interesting and promising branch of research. For example, Liu et al. [33] have proven a rigorous quantum advantage for a classification task engineered from the discrete logarithm problem, although it requires a fault-tolerant quantum computer due to its reliance on a Shor-type data encoding. Beyond that, Refs. [50, 49] formally reveal that supervised QML models are kernel methods. In particular this implies that QKMs can be embedded into the rich mathematical framework of conventional kernel theory [47].
@@
The key idea of kernel methods are kernel functions that implicitly map input data into a higher-dimensional space where the learning problem can often be formulated as a linear model (albeit there is no theoretical guarantee). Quantum kernel methods leverage the principles of quantum mechanics to perform these mappings into the exponentially large Hilbert space of quantum states. In practice, this is realized by parameterized quantum circuits (PQCs) [50, 34, 51, 27], which in the context of quantum kernels are referred to as data encoding circuits. Two common approaches to evaluate the corresponding quantum kernel functions have become established: fidelity quantum kernels (FQKs) [50, 21, 7, 24] and projected quantum kernels (PQKs) [22, 18, 54]. A unified framework for generalized trace-induced quantum kernels that includes FQKs and linear PQKs has been proposed recently in Ref. [18].
@@
Wide-ranging application-oriented research in diverse domains emerged from the promising potentials of QKMs. This ranges from financial classification tasks [37], quantum support vector machines (QSVMs) for modelling tranisition probabilities in health and disability insurance [14], and a quantum kernel classifier for real high-dimensional data from cosmology [42], and QKMs for solving differential equations [40]. Moreover, there are recent efforts to automize the encoding circuit architecture search, cf., e.g., Ref. [43] and references therein.
@@
Despite the variety of previous work, a comprehensive understanding of the wealth of design choices and their underlying mechanisms for various datasets is still incomplete. Our work attempts to make a contribution to closing this gap further by systematically analyzing the diversity of design criteria on a large scale and systematically examining both FQK and PQK approaches. We address this through state-of-the-art hyperparameter search and correlation analysis.
@@
We conduct an extensive benchmarking study for classification and regression tasks using various QKMs such as quantum support vector classification (QSVC), quantum kernel ridge regression (QKRR), and quantum support vector regression (QSVR). Additionally, we generate a broad and general database by exploring nine popular data encoding circuits from the literature, offering a diverse analysis on the impact of the choice of encoding circuit. Beyond that, we propose two different encoding strategies to introduce feature redundancies and systematically compare them. Furthermore, the present work delves deep into the different design aspects of PQKs by examining the effects of measurement operators and outer kernels on model performance. We additionally address the question of which components of the PQK definition are responsible for learning, i.e., the projected quantum circuits versus the outer kernel. Finally, we numerically support all findings by rigorous correlation analysis. In total, this study resulted in over $20,000$ quantum kernel models that were trained to provide an extensive database and to ensure robust and comprehensive insights. To facilitate this, we developed a software tool.
@@
Previous studies have focused mostly on classification tasks. For example, Bowles et al. [8] conducted a large-scale evaluation of popular QML models, including Quantum Neural Networks (QNNs), FQKs and PQKs across various datasets. While QKMs where part of the study, the influence of the different design choices was not a particular focus. Their work was preceded by first attempts to systematically explore certain aspects of quantum model design, cf., e.g., Refs. [38, 28]. Egginger et al. [16] have conducted a hyperparameter study for QKMs which extends the findings of Ref. [53] for PQKs. However, unlike this work, they exclusively focus on a single feature map (Hamiltonian evolution). In contrast to these previous findings, the present study offers an in-depth analysis of QKMs that significantly broadens the scope of investigation in multiple aspects and aims to answer aspects that have been previously untouched. This includes addressing regression problems, encoding mechanisms and a detailed analysis of PQKs. By extending previous findings we aim to contribute to a holistic understanding of QKMs.
@@
This work is organized as follows. Section 2 introduces the theoretical basics of QKMs. Hereafter, we give detailed insights about the study design in Sec. 3, including an overview on the multitude of different quantum kernel models considered in this work, an introduction of the datasets as well as implementation details for the corresponding hyperparameter optimization pipeline and the final experimental setup. In Sec. 4 we first thoroughly investigate model performances and the influence of hyperparameters for classification and regression tasks across all datasets, encoding circuits and quantum kernel models and support corresponding findings with correlation analyses. Secondly, we give an in-depth analysis on different design options within the PQK approach, i.e., choice of outer kernel function and measurement operator. We discuss universal findings and patterns across all experiments and comment on the necessity of entanglement in data encoding circuits in Sec. 5. Finally, we summarize our study in Sec. 6.
2 Theoretical Background
One of the most interesting aspects of QKMs is that they can be formally embedded into the rich and powerful mathematical framework of classical kernel theory [50, 49].
2.1 Conventional Kernel Theory
@@
The key idea behind the conventional kernelized approach to (supervised) machine learning is to find and analyze patterns by transforming the respective learning problem from the original input data domain $\mathcal{X}$ to a higher-dimensional (potentially infinite-dimensional) feature space $\mathcal{F}$, where the learning tasks can often be expressed in a linear form. This mapping is accomplished by a feature map $\phi:\mathcal{X}\to\mathcal{F};\mathbf{x}\mapsto\phi(\mathbf{x})$. Kernels, are real- or complex-valued symmetric and positive semi-definite functions of two input data points, $k:\mathcal{X}\times\mathcal{X}\to\mathbb{C}$, induced by the inner product in the feature space, i.e.,
@@
Hence, less formally, one can think of kernels as similarity measures between two data points $\mathbf{x}$ and $\mathbf{x}^{\prime}$.
A more detailed introduction to conventional kernel theory and respective methods is given in appendices A.1 and A.2.
2.2 Quantum Kernel Methods
In QML we commonly process input data by encoding (embedding) them into quantum states of the form
|$\ket{\psi_{\boldsymbol{\theta}}(\mathbf{x})}$ $=U(\mathbf{x},\boldsymbol{\theta})\ket{0}^{\otimes n}\,,$ (2)
where $\boldsymbol{\theta}$ are variationally trainable parameters to refine the embedding into the quantum Hilbert space $\mathcal{H}^{Q}$. In practice, the unitary encoding operator $U(\mathbf{x},\boldsymbol{\theta})$ can be implemented by a data encoding quantum circuit which manipulates an initial $n$-qubit quantum state |$\ket{0}^{\otimes n}$. This reveals the striking similarity to kernel methods: Both utilize mathematical frameworks that map information into high-dimensional spaces for processing. In particular, as shown in Refs. [50, 49], the central concept of QKMs is that they can be formulated as a classical kernel method (e.g. SVM or KRR) whose kernel is computed using a quantum computer. Since quantum computations inherently feature the quantum mechanical principles of superposition and entanglement, the resulting quantum kernels hold the prospect of designing machine learning models that are able to learn complex problems that are out of reach for conventional machine learning methods [33].
@@ 2.2 Quantum Kernel Methods
Figure 1: Schematic illustration of the basic working principle of QKMs and its two most common approaches to compute respective quantum kernel Gram matrices. Data points are mapped from the input space $\mathcal{X}$ to the quantum Hilbert space $\mathcal{H}^{Q}$ by encoding them into quantum states |$\ket{\psi(\mathbf{x},\boldsymbol{\theta})}$. Access to $\mathcal{H}^{Q}$ is provided by measurements, which can be expressed by inner products of quantum states in full analogy to classical kernel theory. Left: By using the Hilbert-Schmidt inner product and leveraging this fidelity-type metric to define quantum kernels leads to FQKs, cf. Eq. (3). Right: Instead of directly processing quantum states within the quantum Hilbert space it has been shown that it can be beneficial to first project them to an approximate classical representation using, e.g., reduced physical observables. This concept gives rise to the family of PQKs. One of the simplest forms of defining PQKs is given in Eq. (7) and corresponds to measuring $k$-particle reduced density matrices and process the result with a classical kernel function $\kappa$. In both cases (FQK and PQK) the resulting kernel Gram matrices are subsequently passed to a classical kernel algorithm.
To formalize this, we define the density matrix $\rho_{\boldsymbol{\theta}}(\mathbf{x})=\outerproduct{\psi(\mathbf{x},\boldsymbol{\theta})}{\psi(\mathbf{x},\boldsymbol{\theta})}$ | as the corresponding data-encoding feature map [49, 21]. With this, we can leverage the native geometry of the quantum state space and naturally define a quantum kernel using the Hilbert-Schmidt inner product, i.e.,
$k^{\text{FQK}}_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{x}^{\prime})$ $=\mathrm{tr}\left[\rho_{\boldsymbol{\theta}}(\mathbf{x})\rho_{\boldsymbol{\theta}}(\mathbf{x}^{\prime})\right]\,.$ (3)
@@
$k^{\text{FQK}}_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{x}^{\prime})$ $=\left|\Braket{\psi_{\boldsymbol{\theta}}(\mathbf{x})}{\psi_{\boldsymbol{\theta}}(\mathbf{x}^{\prime})}\right|^{2}\,,$ (4)
which, since it represents a fidelity-type metric, is referred to as the fidelity quantum kernel (FQK) [22].
@@
$\rho_{\boldsymbol{\theta}}(\mathbf{x})_{K}$ $=\mathrm{tr}_{j\notin K}\left[\rho_{\boldsymbol{\theta}}(\mathbf{x})\right]\,,$ (5)
@@
$h_{\boldsymbol{\theta}}(\mathbf{x})_{k}$ $=\mathrm{tr}(\rho_{\boldsymbol{\theta}}(\mathbf{x})_{K}O)\,,$
$=\mathrm{tr}(\rho_{\boldsymbol{\theta}}(\mathbf{x})O_{k})\,,$ (6)
where $O_{k}$ represents a $k$-local measurement operator acting on $k\leq n$ qubits, e.g. $P^{\otimes k}\otimes\mathds{1}^{\otimes(n-k)}$, with $P$ a Pauli operator. As such, PQKs can be generally defined as
$k^{\text{PQK}}_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{x}^{\prime})$ $=\kappa\left[h_{\boldsymbol{\theta}}(\mathbf{x})_{k},h_{\boldsymbol{\theta}}(\mathbf{x}^{\prime})_{k}\right]\,.$ (7)
@@
$k^{\text{PQK}}_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{x}^{\prime})$
$=\exp(-\gamma\sum_{k,P}\left[\mathrm{tr}(\rho_{\boldsymbol{\theta}}(\mathbf{x})P_{k})-\mathrm{tr}(\rho_{\boldsymbol{\theta}}(\mathbf{x}^{\prime})P_{k})\right]^{2})$
$=\exp\left[-\gamma F_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{x}^{\prime})\right]\,,$ (8)
where $\gamma\in\mathbb{R_{+}}$ is a hyperparameter. Here, we introduced the latter notation for later use in Sec. 4.4. Note that unless otherwise stated, we are referring to the form as given in Eq. (8) when considering PQKs.
3 Study Design
In this study, we systematically investigate the interplay of various hyperparameters in QKMs with the aim to identify patterns and mechanisms that enhance model performance. A schematic view on the scope of this work is shown in Fig. 2, sketching the models and the basic functionality of the implementation to realize this comprehensive analysis.
The following section details the methodological aspects of our study design.
3.1 Models